# Feasibility Guaranteed Learning-to-Optimize in Wireless Communication Resource Allocation

### v1

@@ Abstract
The emergence of 6G wireless communication enables massive edge device access and supports real-time intelligent services such as the Internet of things (IoT) and vehicle-to-everything (V2X). However, the surge in edge devices connectivity renders wireless resource allocation (RA) tasks as large-scale constrained optimization problems, whereas the stringent real-time requirement poses significant computational challenge for traditional algorithms. To address the challenge, feasibility guaranteed learning-to-optimize (L2O) techniques have recently gained attention. These learning-based methods offer efficient alternatives to conventional solvers by directly learning mappings from system parameters to feasible and near-optimal solutions. This article provide a comprehensive review of L2O model designs and feasibility enforcement techniques and investigates the application of constrained L2O in wireless RA systems and. The paper also presents a case study to benchmark different L2O approaches in weighted sum rate problem, and concludes by identifying key challenges and future research directions.
Index Terms:
resource allocation, learning-to-optimize, feasibility guarantee, deep learning, constrained optimization
@@ I Introduction
The emergence of delay-sensitive 6G applications, such as the internet of things (IoT) and vehicle-to-everything (V2X), transforms wireless resource allocation (RA) into high-dimensional constrained optimization problems that require real-time solutions [1, 2]. While traditional algorithms struggle with high computational cost, learning-to-optimize (L2O) has emerged as a promising alternative, using neural networks (NNs) to significantly accelerate computing process [3]. However, a critical challenge severely hinders the practical deployment of L2O: the lack of feasibility guarantees. Standard NNs may produce solutions that violate the strict operational constraints inherent in wireless systems (e.g., power budgets, rate requirements), potentially compromising system stability and safety. This issue becomes critical in scenarios such as V2X communication, where infeasible solutions compromise system stability and lead to severe safety issues.
To address this feasibility challenge, a common strategy involves post-processing the NN’s raw outputs with traditional projection-based methods, which iteratively map the infeasible solution back into the feasible set. However, this seemingly straightforward approach introduces significant obstacles for L2O frameworks. Such iterative procedures often break the differentiability essential for end-to-end training, complicate gradient computation, and are poorly suited for the GPU-accelerated batch processing that underpins modern deep learning. While some recent works have introduced differentiable convex optimization layers [4], they still face scalability issues and high computational overhead in large-scale applications. Therefore, the fundamental tension between guaranteeing feasibility and maintaining an efficient, end-to-end trainable model has become a primary challenge in L2O.
@@
Recognizing the limitations of simple penalty and projection methods, the L2O research community has developed more sophisticated techniques for various constraints, offering promising avenues for wireless RA. Early approaches, such as [5], incorporated constraint violations as penalty terms in the loss function to guide the training process. To handle equality constraints, OptNet [6] leveraged implicit differentiation via the implicit function theorem to embed feasibility directly into the NN structure. Hybrid approaches such as [7, 8] combined alternative optimization with learning, where primal variables and Lagrange multipliers are alternately updated through dual NN to mitigate constraint violations. Recent studies have shifted toward L2O methods under hard constraints, which align closely with the high reliability demands in V2X and IoT. For instance, [3] introduced a gradient-based correction step to explicitly rectify infeasible outputs based on violated constraints. Homeo-learning [9] and HoP [10] further advanced hard constrained L2O by constructing one-to-one differentiable mappings that guarantee feasibility through invertible neural networks (INNs) and measure transformation, respectively. These approaches offer generalizable and scalable solutions, making them promising candidates for future wireless RA research.
In this article, we focus on leveraging feasibility-guaranteed L2O approaches to address complex constrained optimization problems in wireless RA. We begin by introducing the foundational background of L2O, covering NN architectures and corresponding training strategies. To further illustrate the concept of feasibility guaranteed L2O, we review several representative methods. We then evaluate their effectiveness in practical scenarios through a case study on the quality-of-service (QoS)-aware weighted sum rate (WSR) maximization problem. Finally, we discuss open challenges and outline future research directions in L2O-based wireless RA optimization.
@@ II Fundamentals of Learning to Optimize in Constrained Optimization
The application of constrained L2O in wireless RA typically follows a two-phase framework Fig. 1 which includes optimization problem formulation and obtaining feasible solutions. In the past few decades, various wireless RA problems have been formulated, with objective functions ranging from maximizing sum rates or minimizing total power consumption, while the constraints being the bandwidth or individual rate or power requirements. As illustrated in Fig. 1, the overall process is structured as a modular pipeline. The blue module represents the NN inputs, including system parameters such as channel state information, user requirements, and resource limitations. These inputs define both the optimization objective and the feasibility constraints, as shown in the green shaded region. The orange module corresponds to the learnable component of the L2O framework, responsible for generating initial solutions. Depending on the availability of domain knowledge, this module may adopt either data-driven architectures or model-driven deep learning. Following initial solution generation, the gray shaded module enforces feasibility by projecting the NN outputs to satisfy system constraints. This part encompasses both soft-constrained and hard-constrained methods which is detailed in the next section. The modularity of this design enables flexible adaptation to a wide range of wireless RA problems, particularly in constraint sensitive scenarios. In this section, we focus on the foundational mechanisms of L2O and the initial solution generation phase, including neural architecture design and training strategies.
@@ II-A Overview of L2O Mechanism
Figure 2: Two Perspectives of L2O
@@ II-B NN Designing for Different Application Requirement
Figure 3: Feasibility Guaranteed L2O Methods
@@ II-C Specialization for L2O Training Approaches
III feasibility guaranteed L2O for Constrained Optimization
Although NNs can theoretically approximate optimal input-output mappings, their practical deployment is often hindered by producing infeasible solutions due to a lack of explicit constraint handling. To address this issue, feasibility enforcement mechanisms are integrated into the L2O framework to ensure that generated outputs remain within the feasible region. These mechanisms are generally classified into two categories: soft-constrained approaches, which tolerate minor constraint violations to improve objective performance and computational efficiency; and hard-constrained approaches, which strictly enforce feasibility, potentially at the expense of optimality. This section introduces both approaches along with their implementation details and suitable application scenarios.
@@ III-A Soft-constrained L2O
Penalty-based methods are a widely adopted strategy for soft constraint enforcement in L2O. These methods incorporate constraint violations into the loss function by adding penalty terms that quantify the degree of infeasibility. This allows the model to jointly optimize for the objective and constraint satisfaction during backpropagation. While this approach does not guarantee strict feasibility, it reduces computational overhead and preserves compatibility with diverse neural architectures.
@@
Algorithm-informed architectures embed constraint handling directly into the NN by referring to classical optimization procedures. Rather than treating constraints as post-processing steps or handling them indirectly through penalty terms, these methods embed classical optimization principles, such as the Karush-Kuhn-Tucker (KKT) conditions or the alternating direction method of multipliers (ADMM), into the network design to enforce feasibility. A representative example is the primal–dual learning framework, which employs two coordinated networks: a primal network to optimize the original objective and a dual network to update the Lagrangian multipliers that reduce constraint violation. This approach offers two distinct advantages: First, it establishes a direct correspondence between NN components and optimization variables, enhancing interpretability. Second, it leverages domain knowledge to guide learning, aligning well with the model-driven design principles. However, it also introduces challenges in model training. The coordination between the primal and dual networks is often sensitive to hyperparameter choices such as step sizes and update frequencies. In addition, iterative updates increase memory consumption and may limit scalability in large-scale scenarios
@@ III-B Hard-constrained L2O
This class of methods adopts a two-stage strategy: the NN first generates an initial solution, which is subsequently refined through an iterative correction process to ensure feasibility. A representative example is the DC3 framework [3], which integrates a differentiable correction layer into the L2O pipeline. By computing gradients of constraint violations via automatic differentiation, DC3 iteratively performs gradient updates on the NN output until feasibility is achieved. Another recent approach leverages INNs combined with spherical coordinate transformations [9], where the feasible set is mapped onto a unit hypersphere and projection is performed via binary search along radial directions. The key advantage of iterative methods lies in their reliability: with sufficient correction steps, they can consistently produce feasible outputs regardless of the NN’s initial estimate. However, since these methods primarily focus on constraint satisfaction, they may compromise performance on the original objective, particularly when the correction path diverges from the optimal trajectory.
@@
Structural approaches embed constraint geometry directly into the NN architecture, which enables feasibility and avoid iterative post-corrections. These methods construct explicit bijective mappings between estimated output domains and constraint-satisfying solution domains. One representative technique is the gauge mapping framework [14] which normalizes the output via bounded functions followed by Minkowski projections to maintain geometric alignment with linear constraints. However, such mappings are primarily applicable to bounded linear constraints, such as power or bandwidth. To handle nonlinear or semi-unbounded constraints, our recent work [10] introduces HoP, a polar coordinate-based transformation, where the NN predicts directional vectors and magnitudes that are mapped back to feasible Cartesian coordinates via differentiable trigonometric functions. This design ensures both bijectivity and differentiability, making it particularly suitable for end-to-end integration in L2O frameworks. The primary strength of structural approaches lies in their inherent satisfaction of constraints without relying on penalty terms or iterative refinements, thus avoiding suboptimality and enabling efficient, fully differentiable learning.
@@
Methods / Metrics Optimality Constraints Category Feasibility Training Difficulty Inference Speed
SL [13] High General Weak Easy Fast
Primal-Dual Learning [7] Mid General Weak Mid Fast
Gauge Mapping [14] High Linear Bounded Constraints Hard Mid Mid
Homeo-Learning [9] High Bounded Constraints Hard Hard Mid
DC3 [3] High General Hard Easy Fast
HoP [10] High General Hard Easy Slow
@@ IV Numerical Results and Performance Analysis
This section focuses on applying feasibility-guaranteed L2O approaches to constrained optimization problems in wireless RA. Firstly, we formulate an RA problem based on QoS-aware WSR maximization to benchmark typical L2O approaches. Then, we present a case study comparing the performance of several representative L2O methods under both large-scale and small-scale fading scenarios.
@@ IV-A RA Problem Formulation
The optimization is subject to two primary constraints: (1) a total power budget that includes both transmission and circuit power consumption, and (2) SINR thresholds to ensure QoS for each user. Due to the non-convex objective and constraints, we employ a WMMSE [15] based iterative approach with semidefinite relaxation to derive high-quality feasible solutions that serve as upper-bound baselines.
@@ V Future Opportunities and Open Challenges
While feasibility-guaranteed L2O has been extensively studied in domains such as power systems, operations research, and artificial intelligence, its adoption in wireless communication remains limited where most existing studies in wireless focus on soft-constrained L2O or rely on simple feasibility correction scenarios. This section highlights several open challenges and future research directions for developing constrained L2O methods specific to wireless communication systems.
@@
Wireless communication systems encompass a wide range of constrained optimization problems with diverse and task specific constraints, making feasibility guaranteed L2O a promising solution framework. However, most existing studies in this domain remain confined to simplified settings, typically addressing only basic power constraints through penalty terms or projection sub-routines. In contrast, emerging paradigms such as integrated sensing and communication, reconfigurable intelligent surfaces, and ultra-reliable low-latency communications involve high-dimensional, non-convex, and multi-level constraints that pose substantial challenges to current L2O techniques. Efficiently extending L2O models to handle these various complex constraints will be essential for realizing intelligent and reliable RA in future wireless systems.
@@
Unlike generic optimization problems, RA tasks in wireless systems exhibit distinctive structural properties in both objectives and constraints. For example, beamforming problems often involve quadratic objectives and norm-based power constraints. These inherent features provide opportunities to customized task-specific L2O frameworks via the problem constraint geometry. Rather than aiming for general purpose L2O, such designs can exploit domain knowledge of wireless RA problem to improve model performance and interpretability. For instance, when the feasible set is convex and bounded (e.g., $\ell_{2}$-norm balls), feasibility can be enforced via structured mappings, such as coordinate transformations or adapted activation functions. These task-specific L2O designs not only simplify training but also enhance practicality and performance in deployment.
@@
Most existing L2O approaches are primarily designed for single-variable or decoupled formulations, limiting their effectiveness in capturing the complex interdependencies present in practical wireless systems. In reality, many RA problems, such as multi-user scheduling, beamforming, and coordinated base station control, are inherently multi-variable and highly coupled, where decision variables across users or resources interact in non-separable ways. To accommodate such problems, current methods often reformulate the original formulation into an equivalent single vector representation (e.g., through variable flattening or block-wise stacking). While this simplification enables training within existing L2O frameworks, it obscures the intrinsic coupling structure among variables. Therefore, developing L2O architectures that natively support multi-variable, coupled optimization while preserving problem structure and ensuring feasibility, holds great promise for scalable solutions in multi-variables wireless systems.

### v2

# Feasibility-Aware Learning-to-Optimize for Wireless Resource Allocation

@@ Abstract
The emergence of 6G wireless communication enables massive edge device access and supports real-time intelligent services such as the Internet of things (IoT) and vehicle-to-everything (V2X). However, the surge in edge devices connectivity renders wireless resource allocation (RA) tasks as large-scale constrained optimization problems, whereas the stringent real-time requirement poses significant computational challenge for traditional algorithms. To address the challenge, feasibility-aware learning-to-optimize (L2O) techniques have recently gained attention. These learning-based methods offer efficient alternatives to conventional solvers by directly learning mappings from system parameters to feasible and near-optimal solutions. This article provide a comprehensive review of L2O model designs and feasibility enforcement techniques and investigates the application of constrained L2O in wireless RA systems and. The paper also presents a case study to benchmark different L2O approaches in weighted sum rate problem, and concludes by identifying key challenges and future research directions.
@@ I Introduction
The emergence of delay-sensitive 6G applications, such as the internet of things (IoT) and vehicle-to-everything (V2X), transforms wireless resource allocation (RA) into high-dimensional constrained optimization problems that require real-time solutions [10, 6]. While traditional algorithms struggle with high computational cost, learning-to-optimize (L2O) has emerged as a promising alternative, using neural networks (NNs) to significantly accelerate computing process [2]. However, a critical challenge severely hinders the practical deployment of L2O: the lack of feasibility control. Standard NNs may produce solutions that violate the strict operational constraints inherent in wireless systems (e.g., power budgets, rate requirements), potentially compromising system stability and safety. This issue becomes critical in scenarios such as V2X communication, where infeasible solutions compromise system stability and lead to severe safety issues.
To address this feasibility challenge, a common strategy involves post-processing the NN’s raw outputs with traditional projection-based methods, which iteratively map the infeasible solution back into the feasible set. However, this seemingly straightforward approach introduces significant obstacles for L2O frameworks. Such iterative procedures often break the differentiability essential for end-to-end training, complicate gradient computation, and are poorly suited for the GPU-accelerated batch processing that underpins modern deep learning. While some recent works have introduced differentiable convex optimization layers [13] or differentiable wireless simulation environment [3], they still face scalability issues and high computational overhead in large-scale applications. Therefore, the fundamental tension between feasibility-aware and maintaining an efficient, end-to-end trainable model has become a primary challenge in L2O.
@@
Recognizing the limitations of simple penalty and projection methods, the L2O research community has developed more sophisticated techniques for various constraints, offering promising avenues for wireless RA. Early approaches, such as [14], incorporated constraint violations as penalty terms in the loss function to guide the training process. Hybrid approach, like [9], combined alternative optimization with learning, where primal variables and Lagrange multipliers are alternately updated through dual NN to mitigate constraint violations. Recent studies have shifted toward L2O methods under hard constraints, which align closely with the high reliability demands in V2X and IoT. For instance, [2, 8] introduced a gradient-based correction scheme to explicitly rectify infeasible outputs based on violated constraints. HardNet [7] and HoP [1] further advanced hard constrained L2O by constructing one-to-one differentiable mappings that guarantee feasibility through enforcement layers and measure transformation, respectively. These approaches offer generalizable and scalable solutions, making them promising candidates for future wireless RA research.
In this article, we focus on leveraging feasibility-aware L2O approaches to address complex constrained optimization problems in wireless RA. We begin by introducing the foundational background of L2O, covering NN architectures and corresponding training strategies. To further illustrate the concept of feasibility aware L2O, we review several representative methods. We then evaluate their effectiveness in practical scenarios through a case study on the quality-of-service (QoS)-aware weighted sum rate (WSR) maximization problem. Finally, we discuss open challenges and outline future research directions in L2O-based wireless RA optimization.
@@ II Fundamentals of Learning to Optimize in Constrained Optimization
The application of constrained L2O in wireless RA typically follows a two-phase framework Fig. 1 which includes optimization problem formulation and obtaining feasible solutions. In the past few decades, various wireless RA problems have been formulated, with objective functions ranging from maximizing sum rates or minimizing total power consumption, while the constraints being the bandwidth or individual rate or power requirements. As illustrated in Fig. 1, the overall process is structured as a modular pipeline. The blue module represents the NN inputs, including system parameters such as channel state information, user requirements, and resource limitations. These inputs define both the optimization objective and the feasibility constraints, as shown in the green shaded region. The orange module corresponds to the learnable component of the L2O framework, responsible for generating initial solutions. Depending on the availability of domain knowledge, this module may adopt either data-driven or model-driven deep learning. Following initial solution generation, the gray shaded module enforces feasibility by projecting the NN outputs to satisfy system constraints. This part encompasses both soft-constrained and hard-constrained methods which is detailed in the next section. The modularity of this design enables flexible adaptation to a wide range of wireless RA problems, particularly in constraint sensitive scenarios. In this section, we focus on the foundational mechanisms of L2O and the initial solution generation phase which is the orange module in Fig. 1, including neural architecture design and training strategies.
@@ II-A Overview of L2O Mechanism
Figure 2: Two Perspectives of Feasibility Aware L2O
@@ II-B NN Designing for Different Application Requirement
Figure 3: Feasibility-Aware L2O Methods
@@ II-C Specialization for L2O Training Approaches
III feasibility-Aware L2O for Constrained Optimization
Although NNs can theoretically approximate optimal input-output mappings, their practical deployment is often hindered by producing infeasible solutions due to a lack of explicit constraint handling. To address this issue, feasibility enforcement mechanisms, which are illustrated as the gray module in Fig. 1 and constitute an orthogonal design component independent of neural network architecture choices, are integrated into the L2O framework to ensure that generated outputs remain within the feasible region. These mechanisms can be generally classified into two categories: soft-constrained approaches allow limited constraint violations in exchange for improved objective performance and computational efficiency, which in Fig. 3 include penalty-based and algorithmic methods. In contrast, hard-constrained approaches enforce feasibility strictly throughout the learning and inference processes, potentially sacrificing optimality, and are represented by iterative and structural methods in Fig. 3. This section introduces both approaches along with their implementation details and suitable application scenarios.
@@ III-A Soft-constrained L2O
Penalty-based methods are a widely adopted strategy for soft constraint enforcement in L2O [14]. As shown in penalty part of Fig. 3, these methods incorporate constraint violations into the loss function by adding penalty terms that quantify the degree of infeasibility. This architecture-agnostic approach allows the model to jointly optimize for the objective and constraint satisfaction during backpropagation. While this approach does not guarantee strict feasibility, it reduces computational overhead and preserves compatibility with diverse neural architectures.
@@
Algorithm-informed architectures embed constraint handling directly into the NN by referring to classical optimization procedures. Rather than treating constraints as post-processing steps or handling them indirectly through penalty terms, these methods embed classical optimization principles, such as the Karush-Kuhn-Tucker (KKT) conditions or the alternating direction method of multipliers (ADMM), into the network design to enforce feasibility [9]. A representative example is the primal–dual learning framework, which employs two coordinated networks: a primal network to optimize the original objective and a dual network to update the Lagrangian multipliers that reduce constraint violation. This approach offers two distinct advantages: First, it establishes a direct correspondence between NN components and optimization variables, enhancing interpretability. Second, it leverages domain knowledge to guide learning, aligning well with the model-driven design principles. However, it also introduces challenges in model training. The coordination between the primal and dual networks is often sensitive to hyperparameter choices such as step sizes and update frequencies. In addition, iterative updates increase memory consumption and may limit scalability in large-scale scenarios
@@ III-B Hard-constrained L2O
This class of methods adopts a two-stage strategy: the NN first generates an initial solution, which is subsequently refined through an iterative correction process to ensure feasibility [2, 8]. A representative example is the DC3 framework [2], which integrates a differentiable correction layer into the L2O pipeline where the iterative part of Fig. 3 illustrate the pipline of DC3. By computing gradients of constraint violations via automatic differentiation, DC3 iteratively performs gradient updates on the NN output until feasibility is achieved. Another recent approach constructs efficient and differentiable enforcement layers [7], which guarantees feasibility with input-dependent constraints. The key advantage of iterative methods lies in their reliability: with sufficient correction steps, they can consistently produce feasible outputs regardless of the NN’s initial estimate. However, since these methods primarily focus on constraint satisfaction, they may compromise performance on the original objective, particularly when the correction path diverges from the optimal trajectory.
@@
Structural approaches embed constraint geometry directly into the NN architecture, which enables feasibility and avoid iterative post-corrections. These methods construct explicit bijective mappings between estimated output domains and constraint-satisfying solution domains. To clarify the structural approaches pipeline, Fig. 3 structural part provides the methodology of gauge mapping and HoP. Gauge mapping  [5] which normalizes the output via bounded functions followed by Minkowski projections to maintain geometric alignment with linear constraints. However, such mappings are primarily applicable to bounded linear constraints, such as power or bandwidth. To handle nonlinear or semi-unbounded constraints, recent work [1] introduces HoP, a polar coordinate-based transformation, where the NN predicts directional vectors and magnitudes that are mapped back to feasible Cartesian coordinates via differentiable trigonometric functions. This design ensures both bijectivity and differentiability, making it particularly suitable for end-to-end integration in L2O frameworks. The primary strength of structural approaches lies in their inherent satisfaction of constraints without relying on penalty terms or iterative refinements, thus avoiding suboptimality and enabling efficient, fully differentiable learning.
@@
Methods / Metrics Feasibility Type Optimality Constraints Category Training Difficulty Inference Speed
SL [12] Penalty (soft) High General Easy Fast
Primal-Dual Learning [9] Algorithmic (soft) Mid General Mid Fast
Gauge Mapping [5] Structural (hard) High Linear Bounded Constraints Mid Mid
HardNet [7] Structural (hard) High Bounded Constraints Hard Mid
DC3 [2] Iterative (hard) High General Easy Fast
HoP [1] Structural (hard) High General Easy Slow
@@ IV Numerical Results and Performance Analysis
This section focuses on applying feasibility-aware L2O approaches to constrained optimization problems in wireless RA. Firstly, we formulate an RA problem based on QoS-aware WSR maximization to benchmark typical L2O approaches. Then, we present a case study comparing the performance of several representative L2O methods under both large-scale and small-scale fading scenarios.
@@ IV-A RA Problem Formulation
The optimization is subject to two primary constraints: (1) a total power budget that includes both transmission and circuit power consumption, and (2) SINR thresholds to ensure QoS for each user. Due to the non-convex objective and constraints, we employ a traditional WMMSE [15] based iterative approach with semidefinite relaxation to derive high-quality feasible solutions that serve as upper-bound baselines.
@@ IV-B Performance Evaluation
Moreover, the experimental results indicate that L2O models achieving both strong performance and low violation rates are those combining model-driven neural architectures with feasibility-aware constraint enforcement. This observation suggests that incorporating optimization structure into NN design inherently facilitates a more favorable balance between optimality and feasibility.
@@ V Future Opportunities and Open Challenges
While feasibility-aware L2O has been extensively studied in domains such as power systems, operations research, and artificial intelligence, its adoption in wireless communication remains limited where most existing studies in wireless focus on soft-constrained L2O or rely on simple feasibility correction scenarios. This section highlights several open challenges and future research directions for developing constrained L2O methods specific to wireless communication systems.
@@
Wireless communication systems encompass a wide range of constrained optimization problems with diverse and task specific constraints, making feasibility-aware L2O a promising solution framework. However, most existing studies in this domain remain confined to simplified settings, typically addressing only basic power constraints through penalty terms or projection sub-routines. In contrast, emerging paradigms such as integrated sensing and communication, reconfigurable intelligent surfaces, and ultra-reliable low-latency communications involve high-dimensional, non-convex, and multi-level constraints that pose substantial challenges to current L2O techniques. Efficiently extending L2O models to handle these various complex constraints will be essential for realizing intelligent and reliable RA in future wireless systems.
@@
Unlike generic optimization problems, RA tasks in wireless systems exhibit distinctive structural properties in both objectives and constraints. For example, beamforming problems often involve quadratic objectives and norm-based power constraints. These inherent features provide opportunities to customized task-specific L2O frameworks via the problem constraint geometry. Rather than aiming for general purpose L2O, such designs can exploit domain knowledge of wireless RA problem to improve model performance and interpretability. For instance, when the feasible set is convex and bounded (e.g., $\ell_{2}$-norm balls in QoS problem), feasibility can be enforced via structured mappings, such as coordinate transformations or adapted activation functions. These task-specific L2O designs not only simplify training but also enhance practicality and performance in deployment.
@@
Most existing L2O approaches are primarily designed for single-variable or decoupled formulations, limiting their effectiveness in capturing the complex interdependencies present in practical wireless systems. In reality, many RA problems, such as multi-user scheduling, beamforming, and coordinated base station control, are inherently multi-variable and highly coupled, where decision variables across users or resources interact in non-separable ways. To accommodate such problems, current methods often reformulate the original formulation into an equivalent single vector representation (e.g., through variable flattening or block-wise stacking). While this simplification enables training within existing L2O frameworks, it obscures the intrinsic coupling structure and sensitivity among variables. Therefore, developing L2O architectures that natively support multi-variable optimization while considering variable sensitivity and ensuring feasibility, holds great promise for scalable solutions in multi-variables wireless systems.